Show from national income accounting that

a.

An increase in taxes (while transfers remain constant) must imply a change in net

exports, government purchases, or the saving-investment balance.

b.

An increase in disposable personal income must imply an increase in consumption or an

increase in saving.

c.

An increase in both consumption and saving must imply an increase in disposable

income.

[ For both parts b and c assume there are no interest payments by households or transfer pay

ments to foreigners.] 


1
Expert's answer
2021-08-17T16:56:20-0400

(a)

Taxes are determined by the domestic and foreign trade that is carried out.

We use the following equation;

"TA=G+TR+NX+I-S"

An increase in tax should change the left hand side of the equation.

When transfers remain constant, the change must be higher than the net exports, higher than investment , higher than government spending or lower than savings.

(b)

"YD=S+C"

The conditions of the above equation imply that when disposable income rises, the change that takes place could either be due to an additional saving or an increase in consumption.

(c)

According to the equation "YD=S+C" , it implies that an increase in saving and consumption shows that there is an increase in disposable income.